A bal is thrown horizontally at 9.8m/sec from a tower 40m high.The resultant velocity after 1 sec is....m/sec.

Grade:11

2 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear chakrala,

In a horizontal projectile the horizontal velocity is constant and the vertical velocity changes because of the accelaration due to gravity

Hence Vx = 9.8 m/s

the verticla veloctiy after 1 sec is

Vy = -g*1      (the intial velocity in the verticla direction is zero)

Vy = -9.8 m/s

Hence the resultant velocity after 1 sec is

V = 9.8 m/s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB

vikas askiitian expert
13 years ago

velocity in horizontal direction remains same

so Vx=9.8m/s which is constant

initially velocity in vertical direction is zero

velocity after 1 sec is Vy=Uy+at=0+9.8t            (a=g)

                                         Vy=9.8m/s

 V=9.8i-9.8j in vector form

   =9.8sqrt2 in magnitude